OpenStudy (anonymous):

integration by parts of x^2 * e^x

OpenStudy (anonymous):

let u = x^2 so du = 2x let dv = e^x so v = e^x INT x^2 e^x dx = uv - INT v du = x^2 e^x - INT 2x e^x dx INT 2x e^x dx = 2x e^x - INT 2 e^x dx = 2x e^x - 2e^x finally INT x^2 e^x = x^2 e^x -( 2x e^x - 2 e^x) = e^x(x^2 -2x + 2) + c